Accelerator Update for November 19, 2008 - November 21, 2008
Wednesday Nov 19
The day shift began with Operations monitoring the Tevatron (TeV) store 6574 (127.08E30) and the Recycler stash (240.3E10), with the Antiproton Source stacking (22.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams, with NuMI down due to a horn problem, and with the MiniBooNE experiment taking beam.
Operations terminated store 6574 at 10:32 AM.
Operations prepared CDF and D0 for access at 10:49 AM.
Booster experts reported at 10:54 AM that they were reconnecting a Booster RF station (BRF13) into their RF system.
Operations began preparing Pbar for access at 12:14 PM.
A TeV expert reported at 1:43 PM that he had put the TeV sector A0 kickers into conditioning mode.
A CDF expert reported at 2:27 PM that they had completed their repairs.
Pbar came out of access at 3:02 PM.
Pbar resumed stacking at 3:22 PM.
D0 came out of access at 3:34 PM.
Operations reported at 4:11 PM that the Recycler had stashed antiprotons four times during the day shift.
Operations reported at 4:45 PM that the TeV sector C3 wet engine work was completed.
A TeV power supply expert reported at 7:34 PM that he found and replaced a bad SCR firing card for D0Q3.
Operations reported that they had completed their TeV checkout at 9:15 PM and put the TeV through a dry squeeze.
Operations put the TeV through a wet squeeze at 9:55 PM.
Operations put the TeV into shot setup at 10:46 PM.
Operations reported at midnight that the Recycler had stashed three times during the evening shift.
Thursday Nov 20
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (6.4E10), with Pbar stacking (119.6E10), and with MiniBooNE taking beam.
Operations established store 6578 at 12:14 AM with an initial luminosity of 270.05E30.
At 12:42 AM, the Duty Mechanic reported that the Linac LCW leak was dripping onto a 480-volt breaker panel. The LCW for the Linac Klystrons had to be turned off.
The Duty Electrician reported at 1:17 AM that he had de-energized two Linac breaker panels due to the LCW leak.
The Duty Mechanic reported at 2:08 AM that he had fixed the Linac LCW leak and valved the LCW back on. The Duty Electrician reported that the components of the 480-volt breaker panels did not get wet, but he did dry out water that had collected on the bottom. Linac experts came in and turned things back on. The picture on the right shows the cracked pipe nipple that caused the leak.
Linac experts had Linac RF stations 1-7 up and running by 3:07 AM.
Operations resumed sending beam to Pbar and MiniBooNE at 3:37 AM.
At 4:53 AM, KRF1 tripped off. It reset.
Operations reported at 7:57 AM that the Recycler had stashed four times during the midnight shift.
At 8:12 AM, a Linac power supply (HT4002) tripped off for the first of five times in about 11 minutes. Experts investigated and found the supply was not able to put out enough current.
Linac experts reported at 10:31 AM that they had switched the HT4002 supply with one that was able to deliver the needed current.
A Pbar expert began a study 11:39 AM. The study ended at noon.
Operations reported at 3:58 PM that the Recycler had stashed seven times during the day shift.
At 4:08 PM, store 6578 was lost when the TeV suffered a quench during the crystal collimator studies. The collimator was accidentally moved in the wrong direction.
Operations began preparing the TeV for an access at 6:05 PM to fix the stuck kautzky valve. The picture below shows the valve.
At 6:05 PM, the Booster East Anode Power Supply tripped off due to problems with BRF3 & 5. Operations called in experts.
Operations resumed sending beam to Pbar and MiniBooNE at 6:50 PM.
Cryo system experts completed their Kautzky repair and were out of the TeV by 7:49 PM.
At 8:41 PM, the Cryo Coordinator reported that he was concerned about the TeV sector C4 vacuum. Vacuum technicians were called in.
A Pbar expert began a study at 8:54 PM.
Vacuum experts accessed the TeV at 9:42 PM to install vacuum pumps at C4.
The Pbar study ended at 11:24 PM.
Vacuum experts came out of the TeV at 11:28 PM.
Operations reported at midnight that the Recycler had stashed twice during the evening shift.
Friday Nov 21
The midnight shift began with the TeV recovering from a quench, with Operations monitoring the stash (354.3E10), with no Pbar stack due to their recent studies, and with MiniBooNE taking beam.
Operations put the TeV through a dry squeeze at 12:50 AM.
Operations put the TeV through a wet squeeze at 1:09 AM.
Operations put the TeV into shot setup at 3:20 AM.
Operations established store 6583 at 4:41 AM with an initial luminosity of 318.05E30.
Operations reported at 7:57 AM that the Recycler had stashed three times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today are to put a store into the TeV this afternoon and then prepare MI for an access. The plans for the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 11/21/2008 email Fermilab|
|Security, Privacy, Legal|